# On the Mathematical Foundation of a Decoupled Directional Distortional Hardening Model for Metal Plasticity in the Framework of Rational Thermodynamics

### v2

@@ 5 Discussion
Out of the six material parameters in the model, numerical bounds on only two material constraints, $A_{1}$ and $A_{2}$, can be calculated theoretically, as presented in this paper. The numerical implementation of the proposed model discussed in this study can be performed using the linearized integration technique proposed by Bardet and Choucair [21], which is well suited for rate-type and incremental constitutive equations under generalized loadings. This method transforms nonlinear constitutive behavior and loading constraints into a linear system of ordinary differential equations. A detailed step-by-step description of the algorithm for this model is provided in the Appendix.
The numerical experiments presented in this study serve as a verification of the implementability of the proposed constitutive model. Using the algorithm described in the Appendix, the model was subjected to a uniaxial loading condition with controlled strain increments. The simulation results confirm that the model is numerically stable and capable of capturing key features of elasto-plastic behavior, including yield surface distortion and the evolution of internal variables. However, it must be emphasized that these results, while valuable, do not in themselves validate the predictive capability of the model. The current simulations are performed for demonstration purposes using assumed parameter values. For the model to be used as a reliable predictive tool in material characterization or structural analysis, it must be calibrated against experimental data. Calibration would require the identification of appropriate values for the six material parameters through fitting to, yield surface evolution data.
@@ Acknowledgment
The first author is especially grateful to Professor Y. F. Dafalias, whose pioneering research not only guided this study but also introduced him to the field of rational mechanics.
@@
Appendix: Bardet and Choucair [21] Numerical Procedure
@@
    Initialization: Define initial values for stress $\boldsymbol{\sigma}$, strain $\boldsymbol{\varepsilon}$, internal variables $k$, $\boldsymbol{\alpha}$, and $\mathbf{r}$. If we consider the loading conditions involve either normal stresses ($\sigma_{11}$, $\sigma_{22}$, and $\sigma_{33}$), shear stresses ($\tau_{12}$), or a combination of two normal stresses the stress components $\sigma_{13}$, $\sigma_{23}$, strain components $\varepsilon_{13}$, $\varepsilon_{23}$, and corresponding internal variables are zero. As a result, stress, strain, and internal variables of second rank can be simplified into ($4\times 1$) vectors, and fourth-order tensors as ($4\times 4$) matrices for computational efficiency.
@@
    $\mathbf{S}d\boldsymbol{\sigma}+\mathbf{E}d\boldsymbol{\varepsilon}=(\mathbf{S}+\mathbf{E}\mathbf{B})d\boldsymbol{\sigma}=\mathbf{d}\mathbf{Y},$
@@
    where ${E}$ is Young’s modulus and $\nu$ is the Poisson’s ratio. Using $\mathbf{B}^{trial}$ as $\mathbf{B}$ in the equation at step 2, $d\boldsymbol{\sigma}$ can be calculated.
@@
      $\boldsymbol{\sigma}\leftarrow\boldsymbol{\sigma}+x{d}\boldsymbol{\sigma},\quad\boldsymbol{\varepsilon}\leftarrow\boldsymbol{\varepsilon}+x\mathbf{B}{d}\boldsymbol{\sigma}\vspace{1em}.$
@@
    $\mathbf{B}=\mathbf{B}^{e}+\frac{1}{K_{p}}\frac{\partial f}{\partial\sigma}\otimes\frac{\partial f}{\partial\mathbf{\sigma}}$

### v3

@@ 5 Discussion
Out of the six material parameters in the model, numerical bounds on only two material constraints, $A_{1}$ and $A_{2}$, can be calculated theoretically, as presented in this paper. The numerical implementation of the proposed model discussed in this study can be performed using the linearized integration technique proposed by Bardet and Choucair [2], which is well suited for rate-type and incremental constitutive equations under generalized loadings. This method transforms nonlinear constitutive behavior and loading constraints into a linear system of ordinary differential equations. A detailed step-by-step description of the algorithm for this model is provided in the Appendix A.
The numerical experiments presented in this study serve as a verification of the implementability of the proposed constitutive model. Using the algorithm described in the Appendix A, the model was subjected to a uniaxial loading condition with controlled strain increments. The simulation results confirm that the model is numerically stable and capable of capturing key features of elastoplastic behavior, including yield surface distortion and the evolution of internal variables. However, it must be emphasized that these results, while valuable, do not in themselves validate the predictive capability of the model. The current simulations are performed for demonstration purposes using assumed parameter values. For the model to be used as a reliable predictive tool in material modeling, it must be calibrated against experimental data. Calibration would require the identification of appropriate values for the six material parameters through fitting to, yield surface evolution data.
@@ Acknowledgment
The first and second author is especially grateful to Professor Y. F. Dafalias, whose work not only guided this study but also introduced them to the field of rational mechanics.
@@
Author Contributions
Md Mahmudur Rahman (M.M.R.), Md Mahmudul Hasan Pathik (M.M.H.P.) contributed equally to this work.
M.M.R. : conceptualization, formal analysis, investigation, methodology, writing—original draft, writing—review and editing; M.M.H.P. : conceptualization, formal analysis, investigation, methodology, writing—original draft, writing—review and editing; N.I. : investigation, supervision, writing—original draft, writing—review and editing. All authors gave final approval for publication and agreed to be held accountable for the work performed therein.
Appendix A Appendix: Bardet and Choucair [2] Numerical Procedure
@@
    Initialization: Define initial values for stress $\boldsymbol{\sigma}$, strain $\boldsymbol{\varepsilon}$, internal variables $k$, $\boldsymbol{\alpha}$, and $\mathbf{r}$. If we consider the loading conditions involve either normal stresses ($\sigma_{11}$, $\sigma_{22}$, and $\sigma_{33}$), shear stresses ($\sigma_{12}$), or a combination of two normal stresses the stress components $\sigma_{13}$, $\sigma_{23}$, strain components $\varepsilon_{13}$, $\varepsilon_{23}$, and corresponding internal variables are zero. As a result, stress, strain, and internal variables of second rank can be simplified into ($4\times 1$) vectors, and fourth-order tensors as ($4\times 4$) matrices for computational efficiency.
@@
    $\mathbf{S}d\boldsymbol{\sigma}+\mathbf{E}d\boldsymbol{\varepsilon}=(\mathbf{S}+\mathbf{E}\mathbf{B})d\boldsymbol{\sigma}=\mathbf{d}\mathbf{Y}.$
@@
    where E{E} is Young’s modulus and $\nu$ is the Poisson’s ratio. Using $\mathbf{B}^{trial}$ as $\mathbf{B}$ in the equation at step 2, $d\boldsymbol{\sigma}$ can be calculated.
@@
      $\boldsymbol{\sigma}\leftarrow\boldsymbol{\sigma}+x{d}\boldsymbol{\sigma},\quad\boldsymbol{\varepsilon}\leftarrow\boldsymbol{\varepsilon}+x\mathbf{B}{d}\boldsymbol{\sigma}\vskip 10.00002pt.$
@@
    $\mathbf{B}=\mathbf{B}^{e}+\frac{1}{K_{p}}\frac{\partial f}{\partial\sigma}\otimes\frac{\partial f}{\partial\mathbf{\sigma}}.$